# The sup-norm problem for newforms of large level on $\operatorname{PGL}(n)$

### v1

@@ 1.3.3. Detecting sparse sequences of determinants
However, the unconditional amplifier of [BM15] gives rise to a counting problem where matrices have perfect power determinants, for instance, $n$-th powers. Such a sequence of determinants is too sparse and the method above, averaging over all determinants, produces gross over-counting. Similar issues are well-known already in the classical case $n=2$ (see e.g. the special treatment of square determinants in [HT13]).
@@
for $1\leq\nu\leq n$, where $\chi_{\gamma}$ is the characteristic polynomial of $\gamma$. If this equation is irreducible, then a powerful theorem of Heath-Brown [HB02] provides an adequate non-trivial bound.
To treat the case where the polynomial is reducible, we assume that $n$ is prime to simplify the classification of these degenerate cases. We can thus reduce to counting matrices with $\chi_{\gamma}(X)=(X-m)^{n}$. For $n=2$, this is the special case of parabolic matrices that was also handled in [HT12, Lemma 4.1].
@@
[…]
@@ 2.1. Exterior powers
is an orthonormal basis of $\bigwedge^{n-1}V$, formed by respectively removing each vector $e_{i}$ from the wedge product $e_{1}\wedge\ldots e_{n}$. It is then easy to check that
@@ 2.2. Successive minima
Throughout this paper, we consider successive minima of lattices $L_{z}$ with respect to the unit ball $B^{1}\subset V$ given by the standard inner product. When considering the exterior products of these lattices, successive minima are defined with respect to the compounds of the unit ball, as in the work of Mahler [Mah55] (refer also to [Eve19], Section 3, for a modern treatment).
More precisely, the $k$-th compound of $B^{1}$, denoted here by $B^{k}$, is defined as the convex hull of the points $x_{1}\wedge\cdots\wedge x_{k}$, for all $x_{1},\ldots,x_{k}\in B^{1}$. Mahler notes that $B^{k}$ is a bounded, convex body in $\bigwedge^{k}\mathbb{R}^{n}$, though generally not a sphere (see Section 4 in [Mah55]). Nevertheless, since $B^{k}$ is bounded and 00 is an inner point of $B^{k}$, there are constants $c_{k,n},C_{k,n}>0$ such that
@@
where $B(n,k,r)$ is the ball of radius $r$ inside $\bigwedge^{k}\mathbb{R}^{n}$. As such, the lenght $l$ of the shortest non-zero vector in $\bigwedge^{k}L_{z}$ can be approximated as
@@
A theorem of Mahler (Theorem 3 in [Mah55]; Theorem 3.2 in [Eve19]) relates the successive minima of a lattice to those of its exterior powers. We state here a special case, relevant in this paper.
@@
Let $L$ be a lattice in $\mathbb{R}^{n}$ and let $\lambda_{1},\ldots,\lambda_{n}$ be its successive minima with respect to the unit ball $B^{1}$. Let $\mu_{1}$ be the first successive minimum of the lattice $\bigwedge^{k}L$ with respect to $B^{k}$. Then
@@
We also recall here a classical theorem of Minkowski (see [Cas97, Theorem VIII.1]), stating that
$d(L)\ll_{n}\lambda_{1}\cdots\lambda_{n}\ll_{n}d(L),$ (2.1)
@@
The detailed study of successive minima of $L_{z}$ is crucial in this paper due to the following well-known lemma (see e.g. [BHM16, Lemma 1]), which we apply when counting integral matrices, as explained at the end of Section 3.
@@ 2.3. Iwasawa coordinates and reduction theory
In particular, the statement $z\in\mathbb{H}$ is taken to imply $z\in\operatorname{SL}_{n}(\mathbb{R})$. By the Iwasawa decomposition (see Section 1.2 in [Gol06]), we can take elements in $\mathbb{H}$ to be of the form $z=n(x)\cdot a(y)$, where $n(x)=(x_{ij})_{1\leq i,j\leq n}\in\operatorname{SL}_{n}(\mathbb{R})$ is upper triangular unipotent, meaning that it satisfies
@@
for all $i$, using the Iwasawa coordinates defined above. Reduction theory (see [Bor19, Theorem I.1.4] or [Gol06, Proposition 1.3.2]) shows that
@@
where $w\in\mathbb{H}_{n-1}$ is a matrix in $\operatorname{GL}_{n-1}(\mathbb{R})$. Though not normalised, we can use a variant of the Iwasawa coordinates (it is the one used in Definition 1.2.3 in [Gol06]) to write $w=n(x^{\prime})\cdot a(y^{\prime})$, where
@@
with $h\in\operatorname{SL}_{n-1}(\mathbb{Z})$, acts on $w$ by sending it to $h\cdot w$ and otherwise leaves the last row of $z$ invariant. Reduction theory in degree $n-1$ now implies that there is a parabolic block matrix $g\in\operatorname{SL}_{n}(\mathbb{Z})$ as above so that $g\cdot z=n(x)\cdot a(y)$ with $y_{i}\geq\sqrt{3}/2$ for $i=2,\ldots,n-1$.
More generally one could define a Siegel set $\mathfrak{S}_{\eta}$ for any $\eta>0$ as the set of all $z=n(x)a(y)\in\operatorname{SL}_{n}(\mathbb{R})$ such that $|x_{ij}|\leq 1/2$ and $y_{i}\geq\eta$. The following is a well-known fact in reduction theory, which we state and prove in the version needed in this paper.
Lemma 2.5.
If $z=n(x)a(y)\in\mathfrak{S}_{\eta}$ and $\lambda_{1}\leq\ldots\leq\lambda_{n}$ are the successive minima of $L_{z}$, then
$\lambda_{i}\asymp_{n,\eta}\left\lVert e_{n+1-i}\right\rVert_{z}\asymp_{n,\eta}d_{n+1-i}.$ (2.3)
Proof.
Notice that we can find $n(x^{\prime})\in\operatorname{SL}_{n}(\mathbb{R})$ upper triangular unipotent such that
$z=n(x)a(y)=a(y)n(x^{\prime}).$
One can easily check that
$x^{\prime}_{ij}=x_{ij}\cdot d_{j}/d_{i}=x_{ij}\cdot(y_{n-i}\cdots y_{n-j+1})^{-1}\ll_{\eta}1$
for $i<j$ when $z\in\mathfrak{S}_{\eta}$. Thus, every entry of $n(x^{\prime})$ is bounded uniformly in terms of $\eta$ and so the operator norm of $n(x^{\prime})$ with respect to the Euclidean norm is bounded in terms of $\eta$ and $n$. Since the entries of the inverse $n(x^{\prime})^{-1}$ are polynomials in the $x^{\prime}_{ij}$, we see analogously that its operator norm is also bounded and we can deduce that
$\left\lVert v\right\rVert_{n(x^{\prime})}\asymp_{n,\eta}\left\lVert v\right\rVert$
for all vectors $v\in\mathbb{R}^{n}$. Using coordinates with respect to the standard basis $e_{1},\ldots,e_{n}$, we have
$\left\lVert(c_{1},\ldots,c_{n})\right\rVert_{z}=\left\lVert(c_{1},\ldots,c_{n})\right\rVert_{a(y)n(x^{\prime})}=\left\lVert(d_{1}c_{1},\ldots,d_{n}c_{n})\right\rVert_{n(x^{\prime})}\asymp\left\lVert(d_{1}c_{1},\ldots,d_{n}c_{n})\right\rVert.$
Now $e_{n}\cdot z,\ldots,e_{1}\cdot z$ are linearly independent vectors in $L_{z}$, which implies that $\lambda_{i}\leq\left\lVert e_{n+1-i}\right\rVert_{z}$. Conversely, suppose that $v_{1},\ldots,v_{k}\in L_{z}$ are linearly independent vectors with $\max\left\lVert v_{i}\right\rVert=\lambda_{k}$. In particular, for any $i$ we have $v_{i}=(c_{i1},\ldots,c_{in})\cdot z$ with $c_{ij}\in\mathbb{Z}$ and there is at least one $i\in\{1,\ldots,k\}$ such that $c_{ij}\neq 0$ for some $j\leq n+1-k$ (we are just expressing the fact that $v_{1},\ldots,v_{k}$ cannot be contained in the linear span of the $k-1$ vectors $e_{n+2-k}z,\ldots,e_{n}z$). As such, we have
$\lambda_{k}$ $\geq\left\lVert v_{i}\right\rVert=\left\lVert(c_{i1},\ldots,c_{in})\right\rVert_{z}\asymp_{n,\eta}\left\lVert(d_{1}c_{i1},\ldots,d_{n}c_{in})\right\rVert$
$\geq d_{j}=\frac{d_{j}}{d_{n+1-k}}\cdot d_{n+1-k}\geq\eta^{n+1-k-j}d_{n+1-k}.$
@@
See Lemma 1 in [Ven06]. ∎
Finally, if $\Omega\subset\mathbb{H}$ is a compact set (in particular, it projects to a compact set in the space of lattices $\operatorname{SL}_{n}(\mathbb{Z})\backslash\mathbb{H}$) and $z\in\Omega$, then $\lambda_{1}\gg_{\Omega}1$ by Mahler’s criterion [Bor19, Corollary I.1.9]. The other successive minima must then also be bounded from below, so $\lambda_{i}\gg 1$. By (2.1), we have that
@@
since $z$ has determinant $1$. Thus $\lambda_{2}\asymp 1$ and inductively we find $\lambda_{i}\asymp_{\Omega}1$ for all $i$. We may say $L_{z}$ is an $\Omega$-balanced lattice.
@@ 3. The amplified pre-trace formula
We follow the amplification scheme of Blomer and Maga [BM15], using their archimedean test function but giving also a version that simplifies the sum over Hecke eigenvalues by assuming a conjecture about their sizes.
@@
[…]
@@
[…]
@@ 3.1. The Hecke algebra and Hecke eigenvalues
where $\psi$ is any function on $\Γ\Gamma\backslash\mathbb{H}$. We define the standard Hecke operator as
@@
One computes that the adjoint of $T(p)$ is the operator $T^{\prime}(p)=T_{(1,\ldots,1,0)}(p)$. Let $\lambda(p,\phi_{\varpi})$ be the eigenvalue of $\phi_{\varpi}$ under $T(p)$, so that $\overline{\lambda(p,\phi_{\varpi})}$ is its eigenvalue under $T^{\prime}(p)$. By [BM15, Lemma 4.4] we have
@@
$\Gamma\operatorname{diag}(p,1,\ldots,1)\Gamma\cdot\Gamma\operatorname{diag}(q,\ldots,q,1)\Gamma=\Gamma\operatorname{diag}(pq,q,\ldots,q,1)\Gamma$ (3.1)
in the Hecke algebra, corresponding to the composition $T(p)\cdot T^{\prime}(q)$ (see [BM15, Section 6]).
@@
We use here the normalised eigenvalues $\lambda(p,\varpi)/p^{(n-1)/2}$ as defined in [Gol06, (9.3.5)]. Note that
@@
$\sum_{p\in\mathcal{P}}\frac{|\lambda(p,\varpi_{0})|}{p^{(n-1)/2}}\gg_{\varepsilon}L^{3/4-\varepsilon}.$ (3.2)
@@
The following are standard computations and we refer to Sections 5.1, 5.3, 5.6, 5.7 in [IK04] for more details. Let $\lambda(p)=\lambda(p,\varpi_{0})/p^{(n-1)/2}$ and note that these give the coefficients of the $L$-function attached to $\phi$ or, equivalently, to the automorphic representation $\pi$ generated by $\phi$. Let $L_{\text{RS}}(s)=L(s,\pi\times\tilde{\pi})$ be the Rankin-Selberg $L$-function and define $\Lambda_{\text{RS}}(n)$ to be its coefficients, so that
@@
Then we have $\Lambda_{\text{RS}}(p)=|\lambda(p)|^{2}\log p$. The prime number theorem under GRH states that
$\sum_{n\leq x}\Lambda_{\text{RS}}(n)=x+O_{\varepsilon,\mu}(x^{1/2+\varepsilon}\cdot N^{\varepsilon}).$ (3.3)
@@
for $p\asymp x$. Let $x\gg N^{\delta}$ for some $\delta>0$. The bound above and (3.3) imply that
@@
It is expected that a stronger version of (1.1) holds, that is, with exponent $1$ instead of $3/4$. To prove this we would require the Ramanujan-Petersson conjecture. This would improve the saving in Theorem 2 by doubling the exponent.

### v2

@@ 1.3.3. Detecting sparse sequences of determinants
However, the unconditional amplifier of [7] gives rise to a counting problem where matrices have perfect power determinants, for instance, $n$-th powers. Such a sequence of determinants is too sparse and the method above, averaging over all determinants, produces gross over-counting. Similar issues are well-known already in the classical case $n=2$ (see e.g. the special treatment of square determinants in [22]).
@@
for $1\leq\nu\leq n$, where $\chi_{\gamma}$ is the characteristic polynomial of $\gamma$. If this equation is irreducible, then a powerful theorem of Heath-Brown [23] provides an adequate non-trivial bound.
To treat the case where the polynomial is reducible, we assume that $n$ is prime to simplify the classification of these degenerate cases. We can thus reduce to counting matrices with $\chi_{\gamma}(X)=(X-m)^{n}$. For $n=2$, this is the special case of parabolic matrices that was also handled in [21, Lemma 4.1].
@@
By the Vinogradov notation $f(x)\ll g(x)$ for two functions $f,g$ it is meant that $|f(x)|\leq C\cdot|g(x)|$, at least for large enough $x$, for some $C>0$ called the implied constant. Similarly, for a matrix $X$ and a scalar function $f(X)$ we say that $X=O(f(X))$ when $\left\lVert X\right\rVert\leq C\cdot f(X)$ for some constant $C>0$ and some choice of matrix norm $\left\lVert\cdot\right\rVert$.
@@ 2.1. Exterior powers
is an orthonormal basis of $\bigwedge^{n-1}V$, formed by respectively removing each vector $e_{i}$ from the wedge product $e_{1}\wedge\ldots\wedge e_{n}$. It is then easy to check that
@@ 2.2. Successive minima
Throughout this paper, we consider successive minima of lattices $L_{z}$ with respect to the unit ball $B^{1}\subset V$ given by the standard inner product. When considering the exterior products of these lattices, successive minima are defined with respect to the compounds of the unit ball, as in the work of Mahler [31] (refer also to [18], Section 3, for a modern treatment).
More precisely, the $k$-th compound of $B^{1}$, denoted here by $B^{k}$, is defined as the convex hull of the points $x_{1}\wedge\cdots\wedge x_{k}$, for all $x_{1},\ldots,x_{k}\in B^{1}$. Mahler notes that $B^{k}$ is a bounded, convex body in $\bigwedge^{k}\mathbb{R}^{n}$, though generally not a sphere (see Section 4 in [31]). Nevertheless, since $B^{k}$ is bounded and 0 is an inner point of $B^{k}$, there are constants $c_{k,n},C_{k,n}>0$ such that
@@
where $B(n,k,r)$ is the ball of radius $r$ inside $\bigwedge^{k}\mathbb{R}^{n}$. As such, the length $l$ of the shortest non-zero vector in $\bigwedge^{k}L_{z}$ can be approximated as
@@
A theorem of Mahler (Theorem 3 in [31]; Theorem 3.2 in [18]) relates the successive minima of a lattice to those of its exterior powers. We state here a special case, relevant in this paper.
@@
Let $L$ be a lattice in $\mathbb{R}^{n}$ and let $\lambda_{1}\leq\ldots\leq\lambda_{n}$ be its successive minima with respect to the unit ball $B^{1}$. Let $\mu_{1}$ be the first successive minimum of the lattice $\bigwedge^{k}L$ with respect to $B^{k}$. Then
@@
We also recall here a classical theorem of Minkowski (see [16, Theorem VIII.1]), stating that
(2.1) $d(L)\ll_{n}\lambda_{1}\cdots\lambda_{n}\ll_{n}d(L),$
@@
The detailed study of successive minima of $L_{z}$ is crucial in this paper due to the following well-known lemma (see e.g. [5, Lemma 1]), which we apply when counting integral matrices, as explained at the end of Section 3.
@@ 2.3. Iwasawa coordinates and reduction theory
In particular, the statement $z\in\mathbb{H}$ is taken to imply $z\in\operatorname{SL}_{n}(\mathbb{R})$.
By the Iwasawa decomposition (see Section 1.2 in [20]), we can take elements in $\mathbb{H}$ to be of the form $z=n(x)\cdot a(y)$, where $n(x)=(x_{ij})_{1\leq i,j\leq n}\in\operatorname{SL}_{n}(\mathbb{R})$ is upper triangular unipotent, meaning that it satisfies
@@
Remark 2.4.
When considering the lattice that $z$ defines, the basis given by the Iwasawa decomposition will be particularly useful for computations. In particular, we note that the “last” basis element, or row, has the form
$e_{n}\cdot z=e_{n}\cdot n(x)a(y)=(0,0,\ldots,0,d)$
and the penultimate one is
$e_{n-1}\cdot z=(0,\ldots,0,dy_{1},dx_{n-1,n}).$
Thus, $\left\lVert e_{n}\right\rVert_{z}=d$ and, since there is only one $2\times 2$ minor with non-zero determinant in the matrix formed by the last two rows of $z$,
$\left\lVert e_{n-1}\wedge e_{n}\right\rVert_{z}=d^{2}y_{1}$
Observe here also the useful computation
$\left\lVert e_{2}\wedge\cdots\wedge e_{n}\right\rVert_{z}=\left\lVert e_{1}\right\rVert_{z^{-T}}=(dy_{1}\cdots y_{n-1})^{-1}.$
@@
for all $i$, using the Iwasawa coordinates defined above. [^1] In this paper, a Siegel set more generally refers to a set of points with $x$-coordinates bounded from above in absolute value and $y$-coordinates bounded from below. Reduction theory (see [11, Theorem I.1.4] or [20, Proposition 1.3.2]) shows that
@@
where $w\in\mathbb{H}_{n-1}$ is a matrix in $\operatorname{GL}_{n-1}(\mathbb{R})$. Though not normalised, we can use a variant of the Iwasawa coordinates (it is the one used in Definition 1.2.3 in [20]) to write $w=n(x^{\prime})\cdot a(y^{\prime})$, where
@@
with $h\in\operatorname{SL}_{n-1}(\mathbb{Z})$, acts on $w$ by sending it to $h\cdot w$ and otherwise leaves the last row of $z$ invariant. Reduction theory in degree n1n-1 now implies that there is a parabolic block matrix $g\in\operatorname{SL}_{n}(\mathbb{Z})$ as above so that $g\cdot z=n(x)\cdot a(y)$ with $y_{i}\geq\sqrt{3}/2$ for $i=2,\ldots,n-1$.
@@
See Lemma 1 in [39]. ∎
Finally, if $\Omega\subset\mathbb{H}$ is a compact set (in particular, it projects to a compact set in the space of lattices $\operatorname{SL}_{n}(\mathbb{Z})\backslash\mathbb{H}$) and $z\in\Omega$, then $\lambda_{1}\gg_{\Omega}1$ by Mahler’s criterion [11, Corollary I.1.9]. The other successive minima must then also be bounded from below, so $\lambda_{i}\gg 1$. By (2.1), we have that
@@
since $z$ has determinant $1$. Thus, $\lambda_{2}\asymp 1$ and inductively we find $\lambda_{i}\asymp_{\Omega}1$ for all $i$. We may say $L_{z}$ is an $\Omega$-balanced lattice.
@@
For some parameter $\alpha$, we call a lattice $L$ $\alpha$-balanced if $\lambda_{1}\asymp_{\alpha}d(L)^{1/n}$, where $\lambda_{1}\leq\ldots\leq\lambda_{n}$ are the successive minima of $L$. The lemma above shows that $L_{z}$ is $\Omega$-balanced if $z$ reduces to $\Omega$. Since $\Omega$ will always be fixed, we often call $\Omega$-balanced lattices simply balanced.
Lemma 2.8.
A lattice $L_{z}$ is $\alpha$-balanced if and only if $\lambda_{i}\asymp_{\alpha}d(L)^{1/n}$ for all $i=1,\ldots,n$. Moreover, this property is invariant under scaling and taking duals.
Proof.
This follows from Minkowski’s theorem (2.1), generalising Lemma 2.7. From Lemma 2.1 on the dual lattice and Lemma 2.2 on the successive minima of exterior products, we deduce that $L$ is balanced if and only if the dual $L^{\ast}$ is balanced. ∎
@@ 3. The amplified pre-trace formula
We follow the amplification scheme of Blomer and Maga [7], using their archimedean test function but giving also a version that simplifies the sum over Hecke eigenvalues by assuming a conjecture about their sizes.
@@
$L^{2}(\Gamma_{0}(N)\backslash\mathbb{H})=\int V_{\varpi}\,d\varpi=L^{2}_{\text{cusp}}\oplus L^{2}_{\text{Eis}},$
@@
Recall the Cartan decomposition $G=KAK$, where $A$ is the subgroup of diagonal matrices. The latter has a Lie algebra $\mathfrak{a}$, on which the Weyl group $W$ of $G$ acts. We define the Cartan projection $C(g)\in\mathfrak{a}/W$ of an element $g\in G$ via the Cartan decomposition $g=k_{1}\exp(C(g))k_{2}$, where $k_{1},k_{2}\in K$. Now pick a $W$-invariant norm $\left\lVert\cdot\right\rVert$ on $\mathfrak{a}$. We note that, if $\left\lVert C(g)\right\rVert\ll 1$, then by exponentiating we have
@@ 3.1. The Hecke algebra and Hecke eigenvalues
where $\psi$ is any function on $\Gamma\backslash\mathbb{H}$. We define the standard Hecke operator as
@@
One computes that the adjoint of $T(p)$ is the operator $T^{\prime}(p)=T_{(1,\ldots,1,0)}(p)$. Let $\lambda(p,\phi_{\varpi})$ be the eigenvalue of $\phi_{\varpi}$ under $T(p)$, so that $\overline{\lambda(p,\phi_{\varpi})}$ is its eigenvalue under $T^{\prime}(p)$. By [7, Lemma 4.4] we have
@@
(3.1) $\Gamma\operatorname{diag}(p,1,\ldots,1)\Gamma\cdot\Gamma\operatorname{diag}(q,\ldots,q,1)\Gamma=\Gamma\operatorname{diag}(pq,q,\ldots,q,1)\Gamma$
in the Hecke algebra, corresponding to the composition $T(p)\cdot T^{\prime}(q)$ (see [7, Section 6]).
@@
We use here the normalised eigenvalues $\lambda(p,\varpi)/p^{(n-1)/2}$ as defined in [20, (9.3.5)]. Note that
@@
(3.2) $\sum_{p\in\mathcal{P}}\frac{|\lambda(p,\varpi_{0})|}{p^{(n-1)/2}}\gg_{\varepsilon}L^{3/4-\varepsilon}.$
@@
[…]
@@
Then we have $\Lambda_{\text{RS}}(p)=|\lambda(p)|^{2}\log p$.
The prime number theorem under GRH states that
(3.3) $\sum_{n\leq x}\Lambda_{\text{RS}}(n)=x+O_{\varepsilon,\mu}(x^{1/2+\varepsilon}\cdot N^{\varepsilon}).$
@@
for $p\asymp x$.
Let $x\gg N^{\delta}$ for some $\delta>0$. The bound above and (3.3) imply that
@@
It is expected that a stronger version of (1.3) holds, that is, with exponent 11 instead of $3/4$. To prove this we would require the Ramanujan-Petersson conjecture. This would improve the saving in Theorem 2 by doubling the exponent.